Marissa has $4.30 in quarters and dimes. If she has 6 more quarters than dimes, how many of each coin does she have?
Answer by mananth(16946)

dimes x
quarters( x + 6 )
Convert dollars to cents
Value= 430

10 x 25 x + 150 = 430
35 x = 280
35 x = 280 0
35 x = 280
/ 35
x= 8 dimes = 80 cents
8+6 = 14 quarters
14*25 = 350 cents
350+80 =430 cents